# MathEDU: Towards Adaptive Feedback for Student Mathematical Problem-Solving

### v1

@@ Abstract
Online learning enhances educational accessibility, offering students the flexibility to learn anytime, anywhere. However, a key limitation is the lack of immediate, personalized feedback, particularly in helping students correct errors in math problem-solving. Several studies have investigated the applications of large language models (LLMs) in educational contexts. In this paper, we explore the capabilities of LLMs to assess students’ math problem-solving processes and provide adaptive feedback. The MathEDU dataset is introduced, comprising authentic student solutions annotated with teacher feedback. We evaluate the model’s ability to support personalized learning in two scenarios: one where the model has access to students’ prior answer histories, and another simulating a cold-start context. Experimental results show that the fine-tuned model performs well in identifying correctness. However, the model still faces challenges in generating detailed feedback for pedagogical purposes.
MathEDU: Towards Adaptive Feedback for Student Mathematical Problem-Solving
@@ 1 Introduction
In the post-pandemic era, online learning has emerged as one of the mainstream methods of education (Alqahtani and Rajkhan, 2020; Jafar et al., 2022). Online educational platforms enable students to study anytime, anywhere, offering extensive question banks to assess understanding. However, without immediate teacher support, students may struggle to correct mistakes if they lack a clear grasp of underlying concepts.
Recent advancements in LLMs have prompted numerous studies (Tack and Piech, 2022; Kasneci et al., 2023) to explore their applications in the field of education. Some works (Shen et al., 2021; Yu et al., 2021; Jie et al., 2022) have explored the math word problem-solving capabilities of LLMs, yielding promising results. Leveraging the advanced natural language understanding and generation capabilities of LLMs, we explore their potential in mathematical problem-solving by focusing on the generation of free-text rationalizations.
For effective assessment of student performance, the LLM must possess a robust understanding of mathematical concepts to identify and analyze errors accurately. Additionally, it should provide clear explanations and deliver constructive, adaptive feedback tailored to the student’s reasoning. Yen and Hsu (2023) have employed GPT-3.5 (Ouyang et al., 2022) to simulate both student and teacher roles in order to examine the model’s behavior and grading capabilities. While LLMs can generate comprehensive explanations, findings indicate that they often struggle to accurately interpret students’ mathematical problem-solving processes, leading to incorrect assessments.
To better understand these challenges and explore ways to improve LLMs’ evaluative capabilities, we collect real-world data on student problem-solving processes alongside teacher-written feedback. We extend the existing MathQA (Amini et al., 2019), which contains GRE-level questions requiring advanced mathematical knowledge, by incorporating detailed annotations of students’ problem-solving processes and corresponding teacher feedback. We invited six students to solve these problems, and their answers were reviewed and graded by three mathematics experts. This process resulted in a comprehensive dataset of 4,048 annotated entries. Table 1 provides an example of the annotations, including error types (e.g., incorrect operations), incorrect steps, and teacher feedback addressing misunderstandings. Details of the dataset construction are provided in the following section.
@@
Student Process: $46-36=10$ $10\,\mathrm{km/hr}\times\frac{5}{18}=\frac{100}{36}\,\mathrm{m/s}$ $\frac{100}{36}\times 144=400\,\mathrm{m}$
@@
Rather than focusing solely on whether the final answer matches a predefined solution, this paper investigates how LLMs can evaluate students’ reasoning, recognizing multiple valid approaches. To facilitate this exploration, the task was divided into three subtasks: (1) answer accuracy assessment, (2) error identification, and (3) feedback generation. We investigate the capabilities of LLMs using both direct prompting and fine-tuning approaches. During fine-tuning, we employed three training strategies: single-task training, multi-task training, and end-to-end training. To evaluate the model’s performance in real-world scenarios, experiments were conducted under two conditions: one with access to a student’s prior answer history and another simulating a cold-start scenario without any prior data. In sum, our contributions are threefold: (1) This study investigates the challenge of assessing students’ reasoning processes by exploring the use of LLMs in the context of mathematical education. (2) We present the MathEDU dataset, [^1] The dataset will be released upon acceptance. An anonymized version is available at: https://anonymous.4open.science/r/MathEDU-4628/ designed to address students’ mistakes in math word problem-solving and provide personalized adaptive feedback. (3) We evaluate LLM performance with prompting and LoRA Hu et al. (2021) fine-tuning. Experimental results show that fine-tuning improves accuracy in identifying correctness but remains limited in generating adaptive feedback.
@@ 2.1 Large Language Models for Education
Teachers play a crucial role in identifying and explaining student errors, significantly enhancing understanding and learning efficiency (Robinson and Loeb, 2021; Zhang et al., 2023). To address the lack of interaction and feedback in online learning, researchers have explored automated systems, such as integrating Cognitive Task Analysis (CTA) with LLMs to support student remediation (Graesser et al., 2004; Wang et al., 2024). However, the resource-intensive nature of CTA, requiring expert involvement and complex step design, limits its scalability in routine educational contexts. Some studies leverage LLMs to support instruction by generating distractors for multiple-choice questions (Dave et al., 2021), analyzing students’ answer histories (Gao et al., 2021), and creating personalized instructional materials (He-Yueya et al., 2024). Others use LLMs to simulate students in interactive chats, enhancing classroom simulations for teacher training (Markel et al., 2023). LLMs have also been deployed as virtual tutors, such as generating code explanations in computer science classrooms (MacNeil et al., 2023).
@@ 2.2 Mathematics Education Datasets
Table 2 presents a comparison of existing datasets. The ASSISTments datasets, [^2] https://www.etrialstestbed.org/resources/featured-studies/dataset-papers collected from the free online tutoring platform ASSISTments, focus on addressing the knowledge tracing problem. They consist of grade school math exercises, featuring various question types, including multiple choice, text, and open-ended questions. Multiple versions of the datasets have been released, with data collected during different periods. The Algebra 2005-2006 and Algebra 2006-2007 datasets, collected from the Cognitive Tutors system, capture student responses to Algebra problems.
Dataset #Students #Questions #Answers Solving Process Teacher’s Feedback
ASSISTments 2009 4,217 26,688 346,860
ASSISTments 2009 46,674 179,999 6,123,270
ASSISTments 2015 19,917 - 708,631
ASSISTments 2017 1,709 3,162 942,816
Algebra2005-2006 575 1,084 813,661
Algebra2006-2007 1,840 90,831 2,289,726
Bridge to Algebra 1,146 19,258 3,686,871
EdNet-KT1 784,309 13,169 95,293,926
EdNet-KT2 297,444 13,169 56,360,602
EdNet-KT3 297,915 13,169 89,270,654
EdNet-KT4 297,915 13,169 131,441,538
Junyi 72,630 25,785 16,217,311
Eedi 118,971 27,613 15,867,850
MATHDIAL - 2,861 2,861 $\checkmark$ $\checkmark$
MathEDU 6 4,048 4,048 $\checkmark$ $\checkmark$
Table 2: Comparison of Existing Datasets.
The EdNet dataset (Choi et al., 2020) is a hierarchical dataset designed for educational research. It consists of four levels (KT1 to KT4) that progressively capture more detailed student interactions. These range from question-solving logs to comprehensive action sequences, including behaviors like watching lectures or reading materials. The Junyi dataset (Pojen et al., 2020) consists of student exercise attempt logs collected during 2018-2019. It focuses on high school level Math exercise practice in Taiwan. The Eedi dataset (Wang et al., 2020), collected students’ responses to questions between 2018 and 2020, spans from primary to high school levels. It supports tasks such as predicting student answer correctness and includes additional tasks for assessing question quality and generating personalized question sequences.
Previous datasets for knowledge tracing typically span over a year of diverse student behavioral data, providing valuable insights into student engagement. However, they lack detailed records of students’ step-by-step problem-solving processes in mathematics. Furthermore, they do not include teacher feedback on student errors. To address the limitation, the MATHDIAL dataset (Macina et al., 2023) utilizes LLMs to simulate student responses in mathematical problem-solving, with teachers providing targeted instructional feedback to create educational dialogues.
Although LLMs are capable of simulating student responses, some studies (Aher et al., 2023; Markel et al., 2023) have indicated that these simulated answers are often unrealistic, particularly in maintaining a consistent knowledge level. Consequently, collecting real student problem-solving processes along with the corresponding teacher feedback is essential.
@@ 3 Dataset construction
3.1 Collection of Real-World Student Math Problem-Solving Processes
We construct the MathEDU dataset based on MathQA. [^3] The dataset selection criteria are detailed in Appendix A. We removed the multiple-choice options provided by MathQA, presenting the questions as open-ended tasks. To comprehensively assess students’ mathematical abilities and collect problem-solving processes from individuals with varying skill levels, we recruited six students from different university departments. Before the annotation process, students were asked to attempt 10 questions to ensure their responses adhered to the required format. This step helped avoid issues such as incomplete equations or disorganized layouts, facilitating smoother data collection.
To ensure that each student answered a balanced number of questions across all types (i.e., general, gain, physics, geometry, probability, and other), we sampled 4,500 unique questions from the six types in MathQA, maintaining the proportional distribution of the original dataset. Each student was assigned 750 questions, with the distribution of question types consistent across students. We found that the dataset contained errors such as unrecognizable symbols, unclear problem definitions, and missing accompanying images, which could hinder accurate interpretation. Therefore, students were allowed to note these issues and skip such questions. The missing questions were then supplemented until each student completed 750 questions with problem-solving processes.
Students wrote their problem-solving processes on paper, and their responses were then transcribed into format [^4] https://mathpix.com/ to ensure machine readability. We requested problem-solving steps to be as complete as possible to ensure that the teachers grading their answers could clearly understand their reasoning. If needed, they could also provide verbal explanations to elaborate on their reasoning. For questions they could not solve, students were required to note the reasons. If their handwriting was unclear, we asked them to revise to ensure legibility for scanning and conversion to format. [^5] The annotation guidelines are elaborated in Appendix B.1.  [^6] The question distribution and academic backgrounds of student annotators are detailed in the Appendix B.2.
3.2 Annotation for Error Identification and Teacher Feedback
@@
Each student’s answer was annotated by three experts. In addition to labeling the types of errors and incorrect problem-solving steps, experts provided suggestions for students, including explanations for students’ mistakes and suggestions for correcting them. The questions are reviewed and removed erroneous ones that led to unrecognizable problem-solving processes, resulting in a dataset of 4,048 real student answers. To assess the quality of the annotations, we calculated the inter-rater agreement for error category annotations, using Krippendorff’s Alpha. The result was 0.7818, indicating a substantial level of agreement among the experts in their annotations. When a disagreement occurs, we follow the majority rule to determine the final annotation. In cases where a majority decision could not be reached, the instance was revisited by the three experts to reach a final consensus. Finally, authors with expertise in mathematics education reviewed and refined the labeled data to ensure quality. As a result, our dataset includes 3,050 correct answers and 998 incorrect ones.
4 Data Analysis
4.1 Student Mathematical Abilities
@@
Table 3 presents the correctness of students’ problem-solving processes, highlighting the diverse data we collected even from just six participants. Correctness varied significantly, with the most proficient students achieving up to 87.59%, while those with lower proficiency scored as low as 67.01%. Furthermore, individual strengths differed across areas, with some students reaching 100% accuracy in probability questions, whereas others struggled, achieving only 28.57%.
4.2 Student Problem-Solving Styles
@@
We further analyze variations in students’ problem-solving strategies. Since quantifying these strategies is challenging, we indirectly assess them by evaluating the number of mathematical expressions used by each student. Table 3 shows the average use of equations and written explanations per problem, presented in the last two columns, providing insights into their approaches. This analysis reveals that Students 1, 5, and 6 tend to have more concise problem-solving processes, while Students 2, 3, and 4 have more detailed and complex processes. Notably, Student 2 used an average of 7.3124 equations per problem, the highest among the group.
@@
We also analyzed the average number of English words used, excluding algebraic symbols and expressions, as students could provide verbal explanations to support their solutions. The results reveal significant differences in text usage: Student 6 rarely used words, averaging 0.3106 per problem, while Student 2 provided detailed explanations, averaging 6.7532 words per problem.
5 Methodology
@@
In this section, we focus on how the tutor model is developed to analyze students’ problem-solving processes. The model identifies specific steps where errors occur in incorrect answers and provides targeted feedback. The $i$-th input data consists of the math word question $q_{i}$, the rationale $r_{i}$ provided in MathQA, and the student’s answering process $s_{i}$, which includes both equations and written descriptions of their reasoning. Based on this input, the model performs three main tasks:
@@
Answer Accuracy Assessment: This task determines whether the student’s overall answer is correct based on their problem-solving process, treated as a binary classification with the output label $c_{i}$.
@@
We employ LoRA to fine-tune an LLM. [^8] The details of methods are reported in Appendix D. Three fine-tuning settings are explored. In single-task training, separate models were fine-tuned for each of the three subtasks. In multi-task training, a single model was trained on data annotated for all three subtasks, using distinct prompts to guide the model in performing specific tasks. End-to-end training required the model to sequentially analyze the problem statement and evaluate the student’s answer, providing a complete assessment without dividing the tasks explicitly.

### v2

# MathEDU: Feedback Generation on Problem-Solving Processes for Mathematical Learning Support

@@ Abstract
The increasing reliance on Large Language Models (LLMs) across various domains extends to education, where students progressively use generative AI as a tool for learning. While prior work has examined LLMs’ mathematical ability, their reliability in grading authentic student problem-solving processes and delivering effective feedback remains underexplored. This study introduces MathEDU, a dataset consisting of student problem-solving processes in mathematics and corresponding teacher-written feedback. We systematically evaluate the reliability of various models across three hierarchical tasks: answer correctness classification, error identification, and feedback generation. Experimental results show that fine-tuning strategies effectively improve performance in classifying correctness and locating erroneous steps. However, the generated feedback across models shows a considerable gap from teacher-written feedback. Critically, the generated feedback is often verbose and fails to provide targeted explanations for the student’s underlying misconceptions. This emphasizes the urgent need for trustworthy and pedagogy-aware AI feedback in education.
MathEDU: Feedback Generation on Problem-Solving Processes for Mathematical Learning Support
@@ 1 Introduction
The proliferation of generative AI systems, particularly LLMs, is rapidly reshaping the educational landscape, building on the trends of online learning that became mainstream in the post-pandemic era (Alqahtani and Rajkhan, 2020; Jafar et al., 2022; Tack and Piech, 2022; Kasneci et al., 2023). However, hallucinations pose serious risks by propagating misconceptions or introducing false content, and prior work shows that users are often persuaded by such fluent yet misleading outputs (Mittelstadt et al., 2023).
Figure 1: Example of Model Providing Lengthy but Inappropriate Feedback to the Student’s Solution.
While early research demonstrated the potential of LLMs in solving mathematical word problems (Shen et al., 2021; Yu et al., 2021; Jie et al., 2022), these studies often focused on generating a correct final answer rather than evaluating the student’s reasoning process. This overlooks a critical aspect of learning: understanding how a student arrives at a solution. A significant risk, as summarized by Yen and Hsu (2023), is that LLMs can struggle to correctly interpret students’ mathematical reasoning, sometimes even affirming a student’s flawed logic. This raises a crucial question: How reliably can LLM/large reasoning model (LRM) assess the intermediate steps of a student’s problem-solving process and provide feedback? To investigate this issue, we collect real student problem-solving processes alongside teacher-written feedback. Figure 1 presents a failure case using a real student response, which we evaluated with o1-mini Jaech et al. (2024), an LRM, to determine correctness and generate feedback. We found that the model ignored the student’s actual reasoning process and overwhelmed them with extraneous detail that may increase cognitive load. The generated explanation also failed to clearly pinpoint what was wrong, why it was wrong, or how to fix it.
@@
[…]
@@
To enable broader analysis, this work extends MathQA (Amini et al., 2019) by adding annotations of students’ problem-solving processes together with teacher feedback. In total, 4,048 annotated student solutions were collected, providing a resource for future research. Table 1 provides an example of the annotations, including error types, incorrect steps, and teacher feedback addressing misunderstandings. We investigate how LLMs can evaluate students’ reasoning, recognizing multiple valid approaches. Concretely, we conceptualize the task as three levels: (1) answer correctness classification, (2) error identification, and (3) feedback generation. Answer correctness classification examines whether the model can identify whether the problem-solving process is correct. Error identification assesses whether it can locate the specific error step. Feedback generation is the most advanced level, requiring the model to transform that diagnosis into targeted and pedagogically actionable guidance. In sum, our contributions are threefold: (1) This study investigates the challenge of assessing students’ reasoning processes by exploring the use of LLMs in the context of mathematical education. (2) We present the MathEDU dataset, [^1] https://github.com/NYCU-NLP-Lab/MathEDU designed to address students’ mistakes in math word problem-solving and provide personalized feedback. (3) We evaluate LLM performance with prompting and LoRA Hu et al. (2021) fine-tuning. The results show that fine-tuning enables the model to judge answer correctness more effectively, while generating feedback remains challenging.
@@ 2.1 Large Language Models for Education
Several studies leverage LLMs to support instruction by generating distractors for multiple-choice questions (Dave et al., 2021), analyzing students’ answer histories (Gao et al., 2021), and creating personalized instructional materials (He-Yueya et al., 2024). Others use LLMs to simulate students in interactive chats, enhancing classroom simulations for teacher training (Markel et al., 2023). LLMs have also been deployed as virtual tutors, such as generating code explanations in computer science classrooms (MacNeil et al., 2023). However, LLMs often produce lengthy responses. In learning contexts, these overly detailed responses can cause information overload, reducing students’ focus and diminishing active engagement Lo et al. (2024); Zhang et al. (2025).
@@ 2.2 Mathematics Education Datasets
Several educational datasets have been developed, such as ASSISTments, [^2] https://www.etrialstestbed.org/resources/featured-studies/dataset-papers Algebra 2005–2007, [^3] https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=110444https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=330 EdNet Choi et al. (2020), Junyi Pojen et al. (2020), and Eedi Wang et al. (2020), each capturing large-scale student interaction data across various subjects and grade levels. While these datasets provide valuable records for tasks like knowledge tracing, they primarily log final answers or coarse-grained interaction events, lacking detailed, step-by-step records of students’ problem-solving processes. Teacher-written feedback on student errors is also absent from these studies. MATHDIAL Macina et al. (2023) addresses this gap partially by using LLMs to simulate student responses and pairing them with teacher feedback. Daheim et al. (2024) constructed a dataset combining synthetic and limited annotated student responses to investigate step-level error detection and remediation. Similarly, Gao et al. (2024) presented Math-Minos, leveraging automatically generated stepwise natural language feedback to train verifiers, yet their data relies on model outputs rather than genuine student work. However, prior work Aher et al. (2023); Markel et al. (2023) shows such simulated responses often fail to reflect realistic student reasoning. Thus, collecting real student problem-solving processes together with authentic teacher feedback remains essential for advancing research. MathEDU is built upon authentic student problem-solving processes, each meticulously annotated with teacher-written feedback.
@@ 3 Dataset construction
Question: A train running at the speed of 48 km/hr crosses a pole in 9 seconds. What is the length of the train?
Rationale: Speed = (48 x 5 / 18) m/sec = (40 / 3) m/sec. length of the train = (speed x time). length of the train = (40 / 3 x 9) m = 120 m. answer is c.
Options: a) 140 , b) 130 , c) 120 , d) 170 , e) 160
Correct Option is: C
Table 2: Example of a MathQA Instance.
3.1 Problem-Solving Process Collection
We construct the MathEDU dataset based on MathQA. [^5] The dataset selection criteria are detailed in Appendix A. Table 2 shows an example from MathQA. Each instance contains a math word problem, multiple-choice options, and a rationale, which is the reference solution providing detailed step-by-step reasoning for solving the problem. We removed the multiple-choice options, presenting the questions as open-ended tasks. To comprehensively assess students’ mathematical abilities and collect problem-solving processes from individuals with varying skill levels, we recruited six students from different university departments.  [^6] The annotation guidelines are elaborated in Appendix B.1.777The question distribution and academic backgrounds of student annotators are detailed in Appendix B.2. To ensure that each student answered a balanced number of questions across all types, we sampled 4,500 unique questions from MathQA while preserving the proportional distribution of the original dataset. The six types include “General,” “Gain,” “Physics,” “Geometry (Geo.),” “Probability (Prob.),” and “other.” Each student was assigned 750 questions, with the distribution of question types consistent across students. They wrote the problem-solving processes on paper, and their responses were then transcribed into format [^8] https://mathpix.com/ to ensure machine readability.
Figure 2: Performance of six students across six subject categories (General, Gain, Physics, Geometry, Probability, and Other). Scores are expressed as percentages, with higher values indicating better correctness.
Figure 3: Average numbers of equations and words used by six students in their problem-solving processes.
3.2 Error and Feedback Annotation
@@
Each student’s answer was annotated by three experts. In addition to labeling the types of errors and incorrect problem-solving steps, experts provided suggestions for students, including explanations for students’ mistakes and suggestions for correcting them. The questions are reviewed and removed erroneous ones that led to unrecognizable problem-solving processes, resulting in a dataset of 4,048 real student answers. [^10] During the annotation process, we also asked the experts to flag any cases where the predefined categories did not fit the observed errors. But no such cases were reported. We measured the inter-rater agreement for error category annotations, using Krippendorff’s Alpha. The result was 0.7818, indicating a substantial level of agreement among the experts in their annotations. When a disagreement occurs, we follow the majority rule to determine the final annotation. In cases where a majority decision could not be reached, the instance was revisited by the three experts to reach a final consensus. Finally, we reviewed and refined the labeled data to ensure quality. As a result, our dataset includes 3,050 correct answers and 998 incorrect ones. [^11] The student-solution annotation process took about two months, and the expert annotation took approximately one and a half months.
3.3 Data Analysis
@@
Figure 2 visualizes students’ problem-solving correctness across six subject categories. The radar charts highlight the diversity observed even among the six participants. Correctness varied significantly, with the most proficient students achieving up to 87.59%, while those with lower proficiency scored as low as 67.01%. Furthermore, individual strengths differed across areas, with some students reaching 100% accuracy in probability questions, whereas others struggled, achieving only 28.57%.
@@
We further analyze variations in students’ problem-solving strategies. Since quantifying these strategies is challenging, we indirectly assess them by evaluating the number of mathematical expressions used by each student. Figure 3 shows the average use of equations and written explanations per problem, providing insights into their approaches. This analysis reveals that Students 1, 5, and 6 tend to have more concise problem-solving processes, while Students 2, 3, and 4 have more detailed and complex processes. Notably, Student 2 used an average of 7.3124 equations per problem, the highest among the group.
@@
We also analyzed the average number of English words, excluding algebraic symbols and expressions, as students may provide verbal explanations. The results reveal differences in text usage: Student 6 rarely used words, averaging 0.3106 per problem, while Student 2 provided detailed explanations, averaging 6.7532 words per problem. [^12] More detailed statistics are provided in Appendix D.
4 Grading Problem-Solving Processes
4.1 Task Formulation
@@
In this section, we focus on how the tutor model is developed to analyze students’ problem-solving processes. The $i$-th input data consists of the math word question $q_{i}$, the rationale $r_{i}$ provided in MathQA, and the student’s answering process $s_{i}$, which includes both equations and written descriptions of their reasoning. Based on this input, the model performs three main tasks:
@@
Answer Correctness Classification: This task determines whether the student’s overall answer is correct, treated as a binary classification with the output label $c_{i}$.
@@
4.2 Modeling Approach
Three fine-tuning settings are explored. In single-task training, separate models were fine-tuned for each of the three subtasks. In multi-task training, a single model was trained on data annotated for all three subtasks, using distinct prompts to guide the model in performing specific tasks. End-to-end training required the model to sequentially analyze the problem statement and evaluate the student’s answer, providing a complete assessment without dividing the tasks explicitly.
Specifically, we employ LoRA fine-tuning to train an LLM $\mathcal{M}$, parameterized by $\Phi$. The training data used in LoRA fine-tuning is denoted as $\mathcal{Z}=\{(x_{i},y_{i})\}_{i=1,\ldots,N}$. Given task-specific parameter increment $\Delta\Phi=\Delta\Phi(\Theta)$, we optimize over $\Theta$:
End-to-end without rationale: Similar to the previous approach, this method also uses an end-to-end training approach but excludes the rationale in the input data. The training data used in LoRA fine-tuning is denoted as $\mathcal{Z}=\{(x_{i},y_{i})\}_{i=1,\ldots,N},$ where $x_{i}=\mathcal{P}_{D}(q_{i},p_{i})$ and $y_{i}=(c_{i},E_{i},T_{i})$. Here, $\mathcal{P}_{D}$ represents the prompt that takes $q_{i}$ and $p_{i}$ to generate grading results.
Answer Correctness Classification Error Identification
All General Gain Physics Geo. Prob. Other EM ($\uparrow$) Dis. ($\downarrow$)
Llama3 8B w/o $r$ 66.96% 69.17% 67.23% 63.11% 80.00% 66.67% 56.41% 28.40% 67.61
w/ $r$ 77.04% 77.70% 74.07% 77.50% 84.62% 77.78% 69.77% 11.73% 92.71
Llama3 70B w/o $r$ 88.23% 87.10% 91.50% 86.49% 98.11% 58.82% 93.88% 32.10% 70.23
w/ $r$ 86.65% 83.72% 90.67% 86.06% 98.18% 66.67% 89.80% 29.63% 70.66
GPT-3.5 w/o $r$ 33.83% 28.27% 37.11% 38.37% 45.71% 42.86% 13.79% 30.25% 63.72
w/ $r$ 49.92% 45.79% 53.70% 54.45% 54.05% 30.77% 41.18% 27.78% 60.80
o1-mini w/o $r$ 94.66% 94.23% 94.74% 95.06% 92.00% 94.12% 97.96% 41.98% 56.30
Single-task w/o $r$ 92.39% 90.49% 93.75% 93.66% 96.43% 81.82% 94.12% 24.69% 94.86
w/ $r$ 95.07% 93.01% 96.86% 95.68% 96.43% 90.00% 100.00% 23.46% 96.43
Multi-task w/o $r$ 84.10% 83.15% 80.30% 89.16% 77.27% 71.43% 83.72% 23.46% 94.67
w/ $r$ 94.15% 92.46% 96.05% 94.66% 96.15% 88.89% 95.83% 17.28% 84.99
End-to-end w/o $r$ 92.09% 89.40% 95.71% 94.16% 93.10% 75.00% 94.12% 30.25% 71.01
w/ $r$ 94.38% 92.26% 96.86% 96.09% 96.43% 75.00% 98.04% 40.12% 52.60
Table 3: Results of Answer Correctness Classification (Measured by F-score) and Problem-Solving Error Identification (Measured by Exact Match and Distance) Across Models and Settings.